Telescope.org
This website uses cookies. To read more about our use of cookies, click here.
To not be bothered by this message again, click the OK button.
OK

ART Staff Blog Feed

Dan

Tenerife summer maintenance 2015 update 2

posted by Dan on 13 August 2015

On reviewing last night's alignment run data we found that the quality of the mapped points was insufficient, and would not be usable to create an acceptable pointing model.

Today we have returned to the observatory with the aim of carrying out another alignment run this evening. In preparation for this second alignment run we have adjusted a number of elements of the mount RA drive, in the hope that we will improve the pointing and tracking performance in this axis. We are cautiously optimistic that we have made a breakthrough in this area.. but we won't know for sure until we review the full data set generated by this evening's alignment run.

Dan

Tenerife summer maintenance 2015 update

posted by Dan on 29 July 2015

The summer maintenance trip is now well underway. We have completed the work to re-grease the mount, and to run replacement wiring for the Galaxy camera focuser.

We are currently aiming to bring the system back online shortly with the Galaxy focuser working without its temperature sensor. Once we have the system back online we will be able to continue the work to restore the focuser functionality during the day, minimising the interruption to the operation of the system. To this end we are planning to perform a mapping run tonight with the hope of bringing the system back up tomorrow evening.

Dan

Tenerife summer maintenance 2015: Focuser repair

posted by Dan on 26 July 2015

The delayed 2015 summer maintenance trip has started today. Our focus on this trip will be to repair the faulty focuser on the galaxy camera.

Our initial investigations today have revealed that a fault with the wiring running through the mount has caused the focuser unit to fail. We have run tests with cables bypassing the mount wiring and found that the focus position changes correctly, however even with replacement wiring the temperature sensor is not responding.

We have separately tested the temperature sensor and verified its operation, leading us to believe that there is a fault with the temperature sensor input on the focuser control unit.

We are currently working on a number of solutions to this problem.. but we'll tell you about them later. For now I can say that we're in the process of replacing the faulty wiring and are hoping to re-grease the mount (a part of the regular summer maintenance work) tomorrow. If all goes to plan we will be realigning the system in a few days time, allowing us to return to operation with the cluster and constellation cameras as we work to complete the repairs on the galaxy focuser unit.

Dan

Maintenance Trip June 2014: Mount rebuild timelapse

posted by Dan on 11 June 2014

Each summer we re-grease the Ra and Dec drives on the mount as a part of our routine maintenance. This year we decided the make a timelapse video of the process to show everybody what we get up to on the maintenance trips.
Read more »

Dan

Telescope Operation

posted by Dan on 6 May 2014

Over the last fortnight the operation of the telescope has been somewhat disrupted. This reduction in the capacity of the service has been a result of two unconnected reasons.

Dome code development
The first of the reasons is development of new code to control the dome. Over the last six months we have have made various improvements to the code controlling the dome in an attempt to minimise its impact on images. The dome now attempts to track the mount as it moves in real time through the imaging process. We are hoping to add new code soon which will make the dome aware of the extent to which the dome is obscuring the imaging camera, this will allow the dome to make more intelligent decisions about when it is necessary to rotate to follow the camera.

The changes to the dome code have resulted in a large increase in the amount of communication with the dome, and also have changed the communication from happening in a single thread, to requests coming in from multiple threads. The existing dome control code is thread safe, however the increase in communication appears to now be overwhelming the dome controller; this problem has necessitated a change in the way in which we communicate with the dome. The alterations to the dome control code have required numerous changes throughout the system. Testing and refining of this code has required frequent restarts of the controlserver software. In some circumstances the restart of the controlserver software resets the cool count of the system to zero, this effectively takes the Galaxy camera off-line as it waits to cool again. Whilst this is frustrating, the system does continue to image on the Constellation and Cluster cameras, minimising the amount of lost observing time. I have talked with Chris and believe that we now have a way to indicate on the website when new code development is happening.

Mount pointing errors
We have recently started to have pointing errors with the mount. We are still working to understand the cause of these errors. The system is designed to reject images which have a pointing error in excess of certain predetermined values. Whilst we continue to work on this problem a larger than normal number of the images taken on Galaxy camera will be rejected and discarded by the system. These rejected images are shown in the RTD feed, but their imaging is not recorded in the list of work undertaken in the last night's jobs page; this often makes it looks like the system hasn't done much work in a night.

We have not reached the end of the dome code development cycle, or resolved the mount pointing errors yet. As we move forward with the work we hope to keep you all a bit better informed about what is going on.



Dan

Maintenance Trip Feb 2014 Part 2: More snow.. and a drive around the Observatory

posted by Dan on 24 March 2014

This blog post is going to be a bit of a cheat as I haven't had time to prepare something more substantial. This week I've got some videos for you. On the last trip we filmed the whole of our commute up the mountain in the snowy conditions. I have made a couple of short clips which I hope you all enjoy.
Read more »

Dan

Maintenance Trip Feb 2014 Part 1: Ice, snow and operating the telescope

posted by Dan on 11 March 2014

We have now returned to the UK from our first maintenance trip of 2014. The maintenance trip will be memorable for us for a number of reasons. A lot happened on the trip so instead of a single round up post, I will spread the news over a number of posts in the coming weeks.
Read more »

Dan

BRT Observatory Link Down.. Update

posted by Dan on 13 December 2013

This morning we were greeted with further network failures affecting the IAC weather station and their on site web cameras, however as the day have progressed we have seen a number of systems coming back on-line.
Read more »

Dan

BRT Observatory Link Down..

posted by Dan on 13 December 2013

At roughly 20:30 on the 11th of December we lost communication with the BRT observatory at the Observatorio del Teide, shortly after at around 21:05 we lost communication with our night sky webcam.
Read more »

Dan

Ra trailing update

posted by Dan on 18 September 2013

Trailing in the Ra axis was a noticeable problem with our images in the weeks leading up to the hardware failures in Tenerife. This post talks about the diagnostic work and repairs that were completed to resolve this issue.
Read more »

© 2017 The Open University